OpenStudy (anonymous):

Someone please help explain this to me ! The table shows the outputs y for different inputs x: Input (x) 5 6 7 8 Output (y) 2 5 8 11 Part A: Do the data in this table represent a function? Justify your answer. Part B: Compare the data in the table with the relation f(x) = 2x + 13. Which relation has a greater value when x = 7? Part C: Using the relation in Part B, what is the value of x if f(x) = 75?

OpenStudy (anonymous):

for a, you need to understand the definition of a function. for a function, each input has a single output. for part b, compare the values at x=7 for c, you have\( f(x) = 2x+13\\ 75=f(x)=2x+13\\ 75=2x+13\\ 62=2x\\ x=31\)

OpenStudy (anonymous):

ohhh okay thank you !

OpenStudy (anonymous):

do you need the other 2 more fully explained?

OpenStudy (anonymous):

Only part b how do i compare the values of them? i don't really understand that part

OpenStudy (anonymous):

"Part B: Compare the data in the table with the relation f(x) = 2x + 13. Which relation has a greater value when x = 7? " to find the value at x=7 for f(x), we just plug in x=7 in the relation \(f(7)=2(7)+13=14+13=27\) to find the value at x=7 for the table, just look at the equivalent y value where x=7

OpenStudy (anonymous):

equivalent wasn't the right word for that.